Project

General

Profile

November 5, 2015

Attending: Team IU, Mat, Louise

Action Items:
  • Verify that the operations group gets "watchdog" email
  • Inquire/request ability to edit form names in ECL.
  • Flesh out concept of run cataloging tool.
  • Louise may be requesting a ND HW watch list report soon. Fernanda can provide that with a few days heads up.
Fernanda:
  • Showed new Nearline form. Everyone agreed that it looked good, but we should add instructions to the shifter to describe actions taken to resolve issues.
  • Will show this at Monday ops meeting.
Far Detector:
  • Diblock 4 reported as bad from 6 pm on the 4th though 7 am be good data selection. (Likely do to replace APDs on DiB 4 which ran warm overnight)
  • 7am and 1pm about 80 FEB dropout of readout on dibs 2 and 3 (maintenance at AR). Recovered by shifter
Near Detector:
  • Ran with chiller in stage state over weekended (water too cold)
  • Beam back at 9 am
  • Peak in wrong place ~300us, Fixed by Evan
  • Low intensity running starting on Oct 31